# Data-Driven Prescriptive Analytics Applications: A Comprehensive Survey

### v2

@@ C.4 Domain Expertise (EXP)
Surveyed papers doing uplift analysis for Advertising and Marketing (cf. Appendix B.2) typically used Qini Curve Analysis, a problem-specific technique for determining which customers to target with a treatment after having predicted uplift potential for each [62, 63, 64]. In brief, customers are first sorted by their predicted uplift, and then the percentage of customers to target in this sorted order is plotted on the x-axis of a coordinate system with the cumulative uplift on the y-axis. Assuming an infinite campaign budget, the optimum percentage of customers to target in sorted order can then be found at the maximum point of the curve [62, 64]. To explain this solution, recall that four types of customers are considered in uplift analysis, with campaign treatments for Persuadables and Do-Not-Disturbs respectively causing engagement and disengagement (cf. Table 4). The cumulative uplift function of the Qini curve will not increase unless the campaign targets Persuadables, and will indeed decrease if it targets Do-Not-Disturbs [63,  p. 502].

### v3

@@ C.4 Domain Expertise (EXP)
Surveyed papers doing uplift analysis for Advertising and Marketing (cf. Appendix B.2) typically used Qini Curve Analysis, a problem-specific technique for determining which customers to target with a treatment after having predicted uplift potential for each [62, 63, 64]. In brief, customers are first sorted by their predicted uplift, and then the percentage of customers to target in this sorted order is plotted on the x-axis of a coordinate system with the cumulative uplift on the y-axis. Assuming an infinite campaign budget, the optimum percentage of customers to target in sorted order can then be found at the maximum point of the curve [62, 64]. To explain this solution, recall that four types of customers are considered in uplift analysis, with campaign treatments for Persuadables and Do-Not-Disturbs respectively causing engagement and disengagement (cf. Table 4). The cumulative uplift function of the Qini curve will not increase unless the campaign targets Persuadables, and will indeed decrease if it targets Do-Not-Disturbs [63].